### v1

@@ 1 Introduction
Refer to caption Refer to caption Refer to caption Refer to caption
@@ 2 Related Work
Knowledge Distillation.   Knowledge distillation has proven effective in transferring knowledge from a larger teacher model to a more compact student model Hinton et al. (2015); Buciluǎ et al. (2006); Chen et al. (2020) including distillation from multiple teacher models (You et al., 2017; Yang et al., 2020). Following recent work, we focus on distillation from samples from a model distribution, or symbolic distillation (West et al., 2022), a form of distillation especially common on LLMs, e.g. in instruction tuning (Wang et al., 2023; Taori et al., 2023; Chiang et al., 2023), where instruction-question-answer triples are sourced from a teacher model. In reasoning with LLMs specifically, recent work has distilled reasoning knowledge from a single larger teacher model to a smaller student model Magister et al. (2023); Shridhar et al. (2023); Fu et al. (2023); Ho et al. (2023); Saha et al. (2023); Mukherjee et al. (2023); Mitra et al. (2023); Li et al. (2023); Deng et al. (2023); Liu et al. (2023) using Chain-of-Thought prompting (CoT; Wei et al., 2022) and also, a combination of multiple prompting techniques (Chenglin et al., 2023; Mukherjee et al., 2023; Mitra et al., 2023). Past work has also distilled modular trajectories from a GPT-4 teacher for solving interactive tasks (Chen et al., 2023a; Yin et al., 2023). Overall, different from these single-teacher settings, we delve into the realm of knowledge distillation from multiple teachers. Going one step further, we learn from the interactions between teachers, bringing in fresh challenges on flexibly representing and modeling these interactions.
@@ 3.1 Multi-Agent Interaction Graph (MAG)
Example of a MAG.   MAGs can be generally defined for arbitrary agents and interaction patterns. We focus on reasoning problems as a general class of domains where interaction has positive impacts, defining MAGs for LLM agents on commonsense and math reasoning tasks. Past works have defined several such interaction frameworks Du et al. (2023); Liang et al. (2023); Chen et al. (2023b) for which MAGs can be defined. Of these, we choose ReConcile (Chen et al., 2023b) as our interaction framework because of (1) its performance: it obtains the highest performance across multiple benchmarks, and (2) its flexibility: it allows each agent to converse in natural language following the generic Chain-of-Thought reasoning paradigm and thus, can be readily applied to any downstream task where CoT is applicable. MAGs make minimal assumptions about the contents of nodes and hence, can be similarly defined for agents conversing using other prompting techniques (Chen et al., 2022). Fig. 2 shows an example of a ReConcile-based MAG for a math problem. Since the nodes in a MAG represent model responses (either correct or incorrect), we additionally annotate each node $v\in\mathcal{V}$ with a binary label $y_{v}\in\{0,1\}$, indicating the correctness of the answer. These are marked with green and red circles in Fig. 2. We refer to these two sets of nodes as $\mathcal{V}^{+}$ and $\mathcal{V}^{-}$ respectively such that $\mathcal{V}=\mathcal{V}^{+}\bigcup\mathcal{V}^{-}$. Following the interaction pattern of ReConcile that conditions each subsequent round of interaction on all agent outputs from the previous round, we define edges from all source nodes $v_{i,j}$ to all target nodes $v_{k,j+1}$, i.e., $(v_{i,j},v_{k,j+1})\in\mathcal{E}~{}\forall i,k\in[1,n],\forall j\in[0,r]$.
@@ 3.2 Training Data (MAG) Construction
Benchmarks.   Following the above framework, we construct MAGs for 5 widely-used benchmarks on commonsense and math reasoning: (1) StrategyQA Geva et al. (2021), (2) CommonsenseQA Talmor et al. (2019); Aggarwal et al. (2021), (3) AI2 Reasoning Challenge Clark et al. (2018), (4) GSM8K Cobbe et al. (2021) and (5) MATH Hendrycks et al. (2021). We also experiment with 2 OOD datasets: BoolQ (Clark et al., 2019) and SVAMP (Patel et al., 2021) for which we do not construct MAGs and are exclusively used to test the transfer of our model.
@@ 3.3 Statistics of Training Data (MAGs)
Graph Structure. Lastly, we also show the break-down of different MAG structures. Like nodes, $\mathcal{G}_{0}$ graphs are the most represented in our datasets while $\mathcal{G}_{3}$ graphs are the least, and all graph structures add up to to 1K data points per task. We show examples of MAGs in Appendix D. Using these MAGs as supervision, we train task-specific and multi-task distilled models, as discussed below.
@@ 3.4 MAGDi: Structured Distillation from MAGs
Objective 2: Learning from Negative Reasoning.   Teacher LLMs also make mistakes, particularly when solving more challenging problems. However, instead of discarding these, MAGDi treats them as challenging negatives (generated by a strong teacher) that a student model can learn from by contrasting with positive chains. Given a positive reasoning chain $\textbf{x}^{+}\in\mathcal{V}^{+}$ and a negative reasoning chain $\textbf{x}^{-}\in\mathcal{V}^{-}$, we first extract representations of these chains by performing a weighted average pool over the final layer’s hidden representations of the constituent tokens. We represent these as $h_{\textbf{x}^{+}}\in\mathbb{R}^{d}$ and $h_{\textbf{x}^{-}}\in\mathbb{R}^{d}$ respectively where $d$ is the embedding dimension. Using a projection matrix and a $\tanh$ activation, we further project these embeddings to two scalar scores $s_{\textbf{x}^{+}}\in[-1,1]$ and $s_{\textbf{x}^{-}}\in[-1,1]$. MAGDi then optimizes for the following margin-based objective Cortes & Vapnik (1995) for pairs of positive and negative chains $\textbf{x}^{+},\textbf{x}^{-}$} $\in\mathcal{V}^{+}\times\mathcal{V}^{-}$ in a MAG (see Fig. 3 ‘Learning from Negative Reasoning’).
@@
Objective 3: Learning from Interaction.   Beyond distillation from the correct and incorrect nodes, MAGDi is also intended to distill from the entire conversational structure present in the teachers’ discussion. This would allow the student model to summarize the discussion process and acquire knowledge of how a teacher refines its reasoning chain in each discussion round. Hence, while the previous two objectives assumed nodes to be a disconnected set, MAGDi removes this assumption by also modeling the edges. MAGDi achieves this by augmenting the base student model with a Graph Convolution Network (GCN) module Kipf & Welling (2017). The goal of the GCN is to learn improved, ‘structure-aware’ representations of reasoning chains (nodes) such that the student model learns to discriminate between correct and incorrect nodes in a MAG, eventually leading to better generation of reasoning chains.
@@
where $y_{\textbf{x}}$ is a one-hot encoding of the ground-truth label of a node x.
@@
Table 1: Comparison of structured distillation (MAGDi) with no teacher, single-teacher, and multi-teacher distillation baselines. Firstly, MAGDi outperforms all baselines across all five reasoning benchmarks. On average, MAGDi outperforms the strongest SiT-GPT4 baseline by 4.61% and the no teacher baseline by 10.71%. Secondly, knowledge distillation from each component of MAG improves the student model, as demonstrated by a consistent increase in performance from Level 1 to Level 4.
@@
Multi-Teacher (Ours) MAG MAGDi [Level 4] 74.24 72.56 72.61 52.27 12.76 56.88 [+ 4.61%]
@@ 4 Experimental Setup
We group all methods into three categories of ‘Distillation Source Type’, described as follows. (1) No Teacher.   The lower bound of our distilled models is the zero-shot base model (e.g., Mistral-7B-Instruct). (2) Single-Teacher.   Our next set of baselines only uses training data from a single teacher (out of the three agents used to construct MAGs). This follows multiple prior works (Li et al., 2023; Magister et al., 2023; Fu et al., 2023; Ho et al., 2023) that fine-tune student models on CoT reasoning using the next-token prediction objective (Equation 1). In particular, we fine-tune three student models with one of GPT-4, Bard, or Claude2 as the teacher using only the positive samples from the respective teacher model. We will refer to these Single-Teacher distilled models as SiT. Then SiT-GPT4, for example, will refer to a distilled model trained with only GPT4 as the teacher. (3) Multi-Teacher.   Multi-teacher models correspond to distilled models trained with variants of our proposed MAGDi method. Hence, these also directly demonstrate the utility of the levels in Fig. 2.
@@
5 Results and Analysis
5.1 Main Results
Level 3: Negative reasoning chains further improve distillation.   A student model that additionally learns by contrasting between positive and negative chains improves over all previous levels and the single-teacher baselines. Our MAGDi-AN model obtains an overall average improvement of 3.20% (see MAGDi-AN row versus SiT-GPT4 row).
MAGDi improves inference efficiency compared to ReConcile.   Compared to ReConcile, MAGDi also drastically improves inference efficiency. Because the sizes of gated models like those used in ReConcile are not known, we measure efficiency via number of output tokens generated. [^4] This metric is extremely strict and under-estimates our efficiency gains, as each of the gated LLMs used by ReConcile exceeds 7B parameters, and each example for ReConcile involves multiple inference calls. As shown in Table 2, MAGDi achieves up to a $9$x reduction in token count. MAGDi obtains this efficiency by always answering questions in one inference call; this differs from ReConcile which has to go through up to twelve (expensive) LLM inference calls (involving 3 agents for the initial round and another 3 rounds). Moreover, Figure 4 shows the trade-off between token efficiency and performance (averaged across datasets). While ReConcile (as the upper-bound) has the best performance, it also produces the most tokens (cf. Table 2); the zero-shot and single-teacher models are more efficient than ReConcile but suffer in terms of performance. On the other hand, MAGDi achieves a better balance of efficiency and performance than these baselines, with more efficiency than ReConcile and higher performance than the zero-shot and prior single-teacher distillation methods.
@@ 5.2 Analysis: Generalizability, Scalability, Diversity
MAGDi boosts self-consistency.   We hypothesize that a student model that learns from multiple teachers will exhibit better diversity in its generations. To test this, we combine MAGDi with an ensemble method like Self-Consistency (SC) Wang et al. (2023). SC computes a majority vote over model answers and has proven effective for reasoning tasks; SC’s improvements are predicated on answer diversity (since a majority vote over the versions of the same answer would not yield any improvements). We show that self-consistency with our multi-teacher distilled model outperforms the same with (1) the base model and more importantly, (2) a single-teacher distilled model. As shown in Table 4, on GSM8K, SC between 10 responses improves accuracy by 15% when applied to a MAGDi-trained student, compared to only 4% improvement for the base model (Mistral-7B-Instruct) and 11% for the single-teacher model (SiT-GPT4). Broadly, this suggests that inference-time algorithms relying on a model’s inherent diversity can be boosted from multi-teacher distillation.
6 Conclusion
We have tackled the problem of structured distillation from multi-agent interactions as a way to equip much smaller and more efficient language models with improved reasoning capabilities. To achieve this goal, we proposed a graph-based representation of these interactions, generated graphs for training, and developed a structured distillation method for learning from these interaction graphs. Our results show the effectiveness, generalizability, and scalability of structured distillation across multiple reasoning benchmarks.

### v2

@@ 2 Related Work
Knowledge Distillation.   Knowledge distillation has proven effective in transferring knowledge from a larger teacher model to a more compact student model (Hinton et al., 2015; Buciluǎ et al., 2006; Chen et al., 2020) including distillation from multiple teacher models (You et al., 2017; Yang et al., 2020). Following recent work, we focus on distillation from samples from a model distribution, or symbolic distillation (West et al., 2022), a form of distillation especially common on LLMs, e.g. in instruction tuning (Wang et al., 2023; Taori et al., 2023; Chiang et al., 2023), where instruction-question-answer triples are sourced from a teacher model. In reasoning with LLMs specifically, recent work has distilled reasoning knowledge from a single larger teacher model to a smaller student model (Magister et al., 2023; Shridhar et al., 2023; Fu et al., 2023; Ho et al., 2023; Saha et al., 2023; Mukherjee et al., 2023; Mitra et al., 2023; Li et al., 2023; Deng et al., 2023; Liu et al., 2023) using Chain-of-Thought prompting (CoT; Wei et al., 2022) and also, a combination of multiple prompting techniques (Chenglin et al., 2023; Mukherjee et al., 2023; Mitra et al., 2023). Past work has also distilled modular trajectories from a GPT-4 teacher for solving interactive tasks (Chen et al., 2023a; Yin et al., 2023). Overall, different from these single-teacher settings, we delve into the realm of knowledge distillation from multiple teachers. Going one step further, we learn from the interactions between teachers, bringing in fresh challenges on flexibly representing and modeling these interactions.
@@ 3.1 Multi-Agent Interaction Graph (MAG)
Example of a MAG.   MAGs can be generally defined for arbitrary agents and interaction patterns. We focus on reasoning problems as a general class of domains where interaction has positive impacts, defining MAGs for LLM agents on commonsense and math reasoning tasks. Past works have defined several such interaction frameworks (Du et al., 2023; Liang et al., 2023; Chen et al., 2023b) for which MAGs can be defined. Of these, we choose ReConcile (Chen et al., 2023b) as our interaction framework because of (1) its performance: it obtains the highest performance across multiple benchmarks, and (2) its flexibility: it allows each agent to converse in natural language following the generic Chain-of-Thought reasoning paradigm and thus, can be readily applied to any downstream task where CoT is applicable. MAGs make minimal assumptions about the contents of nodes and hence, can be similarly defined for agents conversing using other prompting techniques (Chen et al., 2022). Fig. 2 shows an example of a ReConcile-based MAG for a math problem. Since the nodes in a MAG represent model responses (either correct or incorrect), we additionally annotate each node $v\in\mathcal{V}$ with a binary label $y_{v}\in\{0,1\}$, indicating the correctness of the answer. These are marked with green and red circles in Fig. 2. We refer to these two sets of nodes as $\mathcal{V}^{+}$ and $\mathcal{V}^{-}$ respectively such that $\mathcal{V}=\mathcal{V}^{+}\bigcup\mathcal{V}^{-}$. Following the interaction pattern of ReConcile that conditions each subsequent round of interaction on all agent outputs from the previous round, we define edges from all source nodes $v_{i,j}$ to all target nodes $v_{k,j+1}$, i.e., $(v_{i,j},v_{k,j+1})\in\mathcal{E}~{}\forall i,k\in[1,n],\forall j\in[0,r]$.
@@ 3.2 Training Data (MAG) Construction
Benchmarks.   Following the above framework, we construct MAGs for 5 widely-used benchmarks on commonsense and math reasoning: (1) StrategyQA (Geva et al., 2021), (2) CommonsenseQA (Talmor et al., 2019; Aggarwal et al., 2021), (3) AI2 Reasoning Challenge (Clark et al., 2018), (4) GSM8K (Cobbe et al., 2021) and (5) MATH (Hendrycks et al., 2021). We also experiment with 2 OOD datasets: BoolQ (Clark et al., 2019) and SVAMP (Patel et al., 2021) for which we do not construct MAGs and are exclusively used to test the transfer of our model.
@@ 3.3 Statistics of Training Data (MAGs)
Graph Structure. Lastly, we also show the break-down of different MAG structures. Like nodes, $\mathcal{G}_{0}$ graphs are the most represented in our datasets while $\mathcal{G}_{3}$ graphs are the least, and all graph structures add up to 1K data points per task. We show examples of MAGs in Appendix D. Using these MAGs as supervision, we train task-specific and multi-task distilled models, as discussed below.
@@ 3.4 MAGDi: Structured Distillation from MAGs
Objective 2: Learning from Negative Reasoning.   Teacher LLMs also make mistakes, particularly when solving more challenging problems. However, instead of discarding these, MAGDi treats them as challenging negatives (generated by a strong teacher) that a student model can learn from by contrasting with positive chains. Given a positive reasoning chain $\textbf{x}^{+}\in\mathcal{V}^{+}$ and a negative reasoning chain $\textbf{x}^{-}\in\mathcal{V}^{-}$, we first extract representations of these chains by performing a weighted average pool over the final layer’s hidden representations of the constituent tokens. We represent these as $h_{\textbf{x}^{+}}\in\mathbb{R}^{d}$ and $h_{\textbf{x}^{-}}\in\mathbb{R}^{d}$ respectively where $d$ is the embedding dimension. Using a projection matrix and a $\tanh$ activation, we further project these embeddings to two scalar scores $s_{\textbf{x}^{+}}\in[-1,1]$ and $s_{\textbf{x}^{-}}\in[-1,1]$. MAGDi then optimizes for the following margin-based objective (Cortes & Vapnik, 1995) for pairs of positive and negative chains $\textbf{x}^{+},\textbf{x}^{-}$} $\in\mathcal{V}^{+}\times\mathcal{V}^{-}$ in a MAG (see Fig. 3 ‘Learning from Negative Reasoning’).
@@
Objective 3: Learning from Interaction.   Beyond distillation from the correct and incorrect nodes, MAGDi is also intended to distill from the entire conversational structure present in the teachers’ discussion. This would allow the student model to summarize the discussion process and acquire knowledge of how a teacher refines its reasoning chain in each discussion round. Hence, while the previous two objectives assumed nodes to be a disconnected set, MAGDi removes this assumption by also modeling the edges. MAGDi achieves this by augmenting the base student model with a Graph Convolution Network (GCN) module (Kipf & Welling, 2017). The goal of the GCN is to learn improved, ‘structure-aware’ representations of reasoning chains (nodes) such that the student model learns to discriminate between correct and incorrect nodes in a MAG, eventually leading to better generation of reasoning chains.
@@
where $y_{\textbf{x}}$ is a one-hot encoding of the label of a node x.
@@
Table 1: Comparison of structured distillation (MAGDi) with no teacher, single-teacher, and multi-teacher distillation baselines. Firstly, MAGDi outperforms all baselines across all five reasoning benchmarks. On average, MAGDi outperforms the strongest SiT-GPT4 baseline by 4.61% and the no teacher baseline by 10.71%. Secondly, knowledge distillation from each component of MAG improves the student model, as demonstrated by a consistent increase in performance from Level 1 to Level 4. Lastly, we also make all nodes available for a multi-teacher baseline, DSS-MT. MAGDi obtains a larger improvement over SiT-GPT4 than DSS-MT does (2.49% vs. 4.61%).
@@
Multi-Teacher MAG MAGDi [Level 4] 74.24 72.56 72.61 52.27 12.76 56.88 [+ 4.61%]
@@ 4 Experimental Setup
We group all methods into three categories of ‘Distillation Source Type’, described as follows. (1) No Teacher.   The lower bound of our distilled models is the zero-shot base model (e.g., Mistral-7B-Instruct). (2) Single-Teacher.   Our next set of baselines only uses training data from a single teacher (out of the three agents used to construct MAGs). This follows multiple prior works (Li et al., 2023; Magister et al., 2023; Fu et al., 2023; Ho et al., 2023) that fine-tune student models on CoT reasoning using the next-token prediction objective (Equation 1). In particular, we fine-tune three student models with one of GPT-4, Bard, or Claude2 as the teacher using only the positive samples from the respective teacher model. We will refer to these Single-Teacher distilled models as SiT. Then SiT-GPT4, for example, will refer to a distilled model trained with only GPT4 as the teacher. (3) Multi-Teacher.   Due to the lack of existing multi-teacher baselines, we first adapt an existing single-teacher method, Distilling Step-by-Step (Hsieh et al., 2023) to a multi-teacher setup. The other multi-teacher baselines (MAGDi-*) correspond to distilled models trained with increasing levels of MAGs; these baselines demonstrate the utility of the levels as defined in Fig. 2.
  • DSS-MT.  Hsieh et al. (2023) propose ‘Distilling Step-by-Step (DSS)’ with a multi-task objective to predict the label and rationale separately. We apply the same multi-task objective to all teachers to directly compare the effectiveness of the DSS objectives on the same MAG data. We refer to this as DSS-MT (DSS with Multi-Teacher).
@@
5 Results and Analysis
5.1 Main Results
Level 3: Negative reasoning chains help.   Additionally, learning by contrasting between positive and negative chains improves over all previous levels. Our MAGDi-AN model obtains an overall average improvement of 3.20% (see MAGDi-AN row versus SiT-GPT4 row).
MAGDi also outperforms adapted single-teacher distillation from prior work.   When using data from all teacher models, MAGDi consistently outperforms DSS-MT. Since the learning source is the same in both methods (i.e., all nodes in MAGs), it underscores the effectiveness of learning from teacher interactions and our proposed objectives.
MAGDi improves inference efficiency compared to ReConcile.   Compared to ReConcile, MAGDi also drastically improves inference efficiency. Because the sizes of gated models like those used in ReConcile are not known, we measure efficiency via number of output tokens generated. [^4] This metric is extremely strict and under-estimates our efficiency gains, as each of the gated LLMs used by ReConcile exceeds 7B parameters, and each example for ReConcile involves multiple inference calls. As shown in Table 2, MAGDi achieves up to a $9$x reduction in token count. MAGDi obtains this efficiency by always answering questions in one inference call; this differs from ReConcile which has to go through up to twelve (expensive) LLM inference calls (involving 3 agents for the initial round and another 3 rounds). Moreover, Figure 4 and Table 10 in Appendix shows the trade-off between token efficiency and performance, averaged across datasets and for each individual dataset respectively. While ReConcile (as the upper-bound) has the best performance, it also produces the most tokens (cf. Table 2); the zero-shot and single-teacher models are more efficient than ReConcile but suffer in terms of performance. On the other hand, MAGDi achieves a better balance of efficiency and performance than these baselines, with more efficiency than ReConcile and higher performance than the zero-shot and prior single-teacher distillation methods.
@@ 5.1 Main Results
MAGDi effectively transfers multi-agent capabilities into a single student.   In Table 3, we compare distilled student models to their teachers by reporting results for (1) GPT4, as the upper-bound performance of a single-teacher, (2) ReConcile, as the upper-bound of multi-agent teacher (using GPT4, Claude2, Bard), (3) SiT-GPT4, as distillation from the single-teacher, and (4) MAGDi as distillation from multi-agent teacher. We also report the relative improvement from ‘single’ to ‘multi’, both with and without distillation. The relative improvement from SiT-GPT4 to MAGDi (12.70%) is much higher than that from GPT-4 to the multi-teacher system ReConcile (5.79%), highlighting the effectiveness of MAGDi in transferring multi-agent capabilities into a single student model.
@@ 5.2 Analysis: Generalizability, Scalability, Diversity
Table 3: Comparison of single-teacher and multi-teacher distilled students with their respective teachers. The relative improvement from SiT-GPT4 to MAGDi (12.70%) is higher than that from GPT-4 to the multi-teacher system ReConcile (5.79%).
StrategyQA CSQA ARC-c GSM8K MATH Avg
GPT4 75.60 73.30 94.50 90.70 39.00 74.62
ReConcile 87.70 78.70 96.30 92.10 41.00 79.16
% improved 13.80 6.86 1.87 1.52 4.88 5.79
SiT-GPT4 69.96 66.87 68.91 47.38 8.24 52.27
MAGDi 74.24 72.56 72.61 52.27 12.76 56.88
% improved 5.77 7.84 5.10 9.36 35.42 12.70
@@
MAGDi boosts self-consistency.   We hypothesize that a student model that learns from multiple teachers will exhibit better diversity in its generations. To test this, we combine MAGDi with an ensemble method like Self-Consistency (SC) (Wang et al., 2023). SC computes a majority vote over model answers and has proven effective for reasoning tasks; SC’s improvements are predicated on answer diversity (since a majority vote over the versions of the same answer would not yield any improvements). We show that self-consistency with our multi-teacher distilled model outperforms the same with (1) the base model and more importantly, (2) a single-teacher distilled model. As shown in Table 5, on GSM8K, SC between 10 responses improves accuracy by 15% when applied to a MAGDi-trained student, compared to only 4% improvement for the base model (Mistral-7B-Instruct) and 11% for the single-teacher model (SiT-GPT4). Broadly, this suggests that inference-time algorithms relying on a model’s inherent diversity can be boosted from multi-teacher distillation.
GCN outperforms self-attention for modeling MAGs.   To demonstrate the effectiveness of the GCN in encoding MAGs, we compare it to two alternative approaches. These remove the explicitly-defined graph structure of the GCN (which is determined by the MAG) and instead aim to automatically learn the interactions with multi-head attention layers (Vaswani et al., 2017). Token-level attention: Here we linearize the entire interaction (across agents and rounds) into a single sequence of tokens, so that any token (across all reasoning chains, rounds, and agents) can attend to any other token. Node-level attention: Here we keep the nodes intact, i.e., each node still represents a whole reasoning chain, but we linearize them into a sequence of nodes so that any node can attend to any other node. Results presented in Table LABEL:tab:structure_ablation show that visualizing the entire interaction as a sequence of tokens results in long token lengths and removes the boundaries between the nodes (reasoning chains), leading to significantly worse performance (first row). Node-level attention improves slightly over token-level attention by keeping the nodes (chains) intact and modeling the interactions between them with self-attention layers (second row). However, our GCN-based approach performs the best across all datasets. While the transformer could theoretically model any graph structure, it may also require a large amount of training data and tuning to infer the correct interaction graph structure. Hence, we find that defining the interaction patterns with MAGs apriori and modeling them with a GCN is more performant and data-efficient. Note that MAGDi’s main contributions are in defining structured representations of multi-agent interactions and then distilling such interaction data into a weaker student model. This contribution is independent of the exact graph representation learning module, and going forward, we hope that our results will motivate newer methods of modeling these multi-agent interactions.
@@
Table 6: Comparison of different graph modeling methods. MAGDi with GCN consistently outperforms attention-based variants for modeling interaction graphs.
Method StrategyQA CSQA ARC-c GSM8K MATH
Token Attn 68.56 65.04 70.31 51.17 8.86
Node Attn 69.87 71.16 70.01 51.63 9.22
GCN 74.24 72.56 72.61 52.27 12.76
6 Discussion and Conclusion
We have tackled the problem of structured distillation from multi-agent interactions as a way to equip much smaller and more efficient language models with improved reasoning capabilities. To achieve this goal, we proposed a graph-based representation of these interactions, generated graphs for training, and developed a structured distillation method for learning from these interaction graphs. Our results showed the effectiveness, generalizability, and scalability of structured distillation across multiple reasoning benchmarks.
While MAGDi relies on interactions between LLMs for supervision, its modular design makes it well-suited even to scenarios where such data may be limited. Revisiting MAGDi’s four objectives, we split its use cases into the following four categories. (1) No teacher data is available: When no teacher data is available, we demonstrate MAGDi’s ability to generalize in a zero-shot manner (Table 4). (2) Only single-teacher data is available: When supervision from only one teacher model is available, we can leverage the first objective in MAGDi (Level 1). (3) Having multi-teacher data but no interaction: When multiple teacher sources are present but the interactions are absent, we show gains from the first two MAGDi objectives (Level 3). (4) Having multi-teacher interaction data: This is when the full MAGDi model (Level 4) can be applied to maximize performance. Thus, our four levels succinctly depict the adaptability of MAGDi based on the amount and nature of data availability, providing a structured framework for implementing its objectives across varying data scenarios and applications.